If X-double-bar = 23 ounces, σ = 0.4 ounces, and n = 16, what will be the ±3σ control limits for the x-bar chart?

   

22.70 to 23.30 ounces

   

21.8 to 24.2 ounces

   

22.25 to 23.75 ounces

   

23 ounces

which is the correct answer?
Answer #1

Process average=23

σ = 0.4 ounces

σx=σ/sqrt(n) where n= sample size

σx=0.4/sqrt(16)=0.4/4=0.1

UCL=Process average+zσx   LCL=Process average-zσx

UCL=23+3*0.1=23.3

LCL=23-3*0.1=22.7